OpenStudy (anonymous):

what is the derivative of 3√4x^5

OpenStudy (anonymous):

\[3(\sqrt{4x})^{5}\] or \[3\sqrt{4x ^{5}}\]

OpenStudy (anonymous):

the second answer

OpenStudy (anonymous):

Ok, so you can rewrite that \[6\sqrt{x ^{5}}\] And I would rewrite that to \[6(x^{5})^{\frac{ 1 }{ 2}}\] Now use the chain rule. Do you know how to do that?

OpenStudy (anonymous):

ok thank you for your help.

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.